If I am chasing someone and step for step I am right behind I have no chance of coming up even with him. So why is it that if I run at a forty-five degree angle from his angle of direction, I am eventually able to come up even with him, even though at a distance from him? We walked the distance and it works in practice.
Answer by stanbon(75887)

If I am chasing someone and step for step I am right behind I have no chance of coming up even with him. So why is it that if I run at a forty-five degree angle from his angle of direction, I am eventually able to come up even with him, even though at a distance from him? We walked the distance and it works in practice.
-----
If he walks 10 feet you will walk 10(sqrt(3)/2) = 10*0.86 = 8.6 ft.
==================
You will not be "even with him".
Cheers,
Stan H.